OpenStudy (anonymous):

write an example of an algebraic expressions that always has the same value regardless of the value of the variable. Need help with hw please

jimthompson5910 (jim_thompson5910):

here's a simple example: if you start with any number x and you subtract x from it you get x - x = 0x = 0 so regardless of the value of x, the value of x-x will be 0

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

or another example: if you start with 78x and you divide it by 13x, you get 78x/13x = (78/13)*(x/x) = (6)*(1) = 6 therefore, the expression 78x/13x is always equal to 6 regardless of the value of x (note: x can't be zero)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

another example 5x + 10 = 7x - 22

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u r welcome